OpenStudy (anonymous):

f(x)=1/x^2-3 substitute h for x in function to solve for f(h)

OpenStudy (zzr0ck3r):

put a h everywhere you see a x

OpenStudy (anonymous):

so, f(h)=1/h(x^2-3)?

OpenStudy (anonymous):

or would it be f(h)=1/h^2-3 without the parentheses? @zzr0ck3r

OpenStudy (zzr0ck3r):

f(h) = 1/(h^2-3)

OpenStudy (zzr0ck3r):

is x^2-3 on the bottom or just x^2?

OpenStudy (anonymous):

It's x^2-3